A hydrogen atom is in the n = 6 state. Determine, according to quantum mechanics, (a) the total energy (in eV) of the atom, (b) the magnitude of the maximum angular momentum the electron can have in this state, and (c) the maximum value that the z component Lz of the angular momentum can have.
Answer #1

$A hydrogen atom is in $ 6^{th}$ state , so according to quantum mechanics $\\a.$ The energy of the atom is $\epsilon =-\frac{13.6eV}{6^2}=0.377eV.\\b.$ The maximum value of l=n-1=5, so the maximum value of angular momentum L=$\sqrt{l(l+1))}\hbar=\sqrt{30} \hbar.\\c.$ The maximum value of $m_l=5 $ then maximum value of z-component of the angular momentum is $L_z=m_l\hbar=5\hbar